OpenStudy (anonymous):

How do I solve the problem tanx=1/2 using the double angle formula to find sin2x, tan2x and cos 2x

myininaya (myininaya):

build a right triangle for tan(x)=1/2 recall tan(x)=opp/adj find the hyp given the opp and the adj side you will have to use the double angle identities on sin(2x) and cos(2x)